# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

### v1

@@ 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in (6) means
${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
(15) can be computed element-wise as follows (see Appendix C for a derivation):
@@
Interpretation: Soft-thresholding in (16) to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$. Given that graph Laplacians—second difference matrices on graphs—are high-pass filters [3], we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank [30], where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$. Using (16) for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved. For example, the constant (low-frequency) signal ${\mathbf{1}}$ passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$. Thus, attenuation of the high-pass channel means (16) is a low-pass filter.
@@ 3.4.3 Updating Lagrange Multipliers
For objective (6), we follow standard ADMM procedure to update Lagrange multiplier ${\boldsymbol{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
${\boldsymbol{\gamma}}^{\tau+1}$ $={\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\boldsymbol{\gamma}}_{u}^{\tau+1}$ and ${\boldsymbol{\gamma}}_{d}^{\tau+1}$ in objective (9) must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~{}~{}~{}~{}~{}~{}~{}~{}~{}{\boldsymbol{\gamma}}_{d}^{\tau+1}={\boldsymbol{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
@@ 4 Unrolled Neural Network
We unroll our iterative ADMM algorithm to an interpretable neural net. The crux to our unrolled neural net is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the self-attention mechanism in classical transformers [19].
@@ 4.1 ADMM Algorithm Unrolling
We implement each iteration of our algorithm in Section 3.4 as a neural layer. Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems (10) for ${\mathbf{x}}^{\tau+1}$, (11) for ${\mathbf{z}}_{u}^{\tau+1}$, and (12) for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via (16). CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix D for detailed implementation). Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer. Multipliers ${\boldsymbol{\gamma}}^{\tau+1},{\boldsymbol{\gamma}}_{u}^{\tau+1},{\boldsymbol{\gamma}}_{d}^{\tau+1}$ are then updated via (17) and (18) to complete one layer. Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\boldsymbol{\Theta}}_{b}$. See Fig. 3 for an illustration.
@@ 4.2 Self-Attention Operator in Transformer
We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation [19]. Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$. The affinity between tokens $i$ and $j$, $e(i,j)$, is the dot product of linear-transformed ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{Q}},{\mathbf{K}}\in\mathbb{R}^{E\times E}$ are the query and key matrices, respectively. Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
@@ 4.3 Graph Learning Modules
Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$. For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng and Tassiulas [9], and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix E.3 for details). As done in Thuc et al. [18], the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
@@
where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$. Distance in (21) is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$. Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
@@
Remark: Interpreting distance $\text{d}^{u}(i,j)$ as $-e(i,j)$, edge weights $w^{u}_{i,j}$’s in (22) are essentially attention weights $a_{i,j}$’s in (19), and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix ${\mathbf{M}}$ can be much smaller than large and dense query and key matrices, ${\mathbf{Q}}$ and ${\mathbf{K}}$. Further, there is no value matrix ${\mathbf{V}}$ required, since the output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—(10), (11), (12) and (16)—derived from our optimization of objective (4).
Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$. For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$. Like (21), we compute Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$, given feature vectors ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$, using a learned PSD metric matrix ${\mathbf{P}}$. Weight $w_{j,i}^{d}$ of directed edge $[i,j]\in{\mathcal{E}}^{d}$ is then computed using exponential functions and normalization, as done in (22) for an undirected edge.
Parameters learned via back-propagation by the two graph learning modules at block $b$—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, ${\mathbf{M}}$ and ${\mathbf{P}}$—are denoted by ${\boldsymbol{\Phi}}_{b}$.
@@ 5.1 Experimental Setup
We evaluate our model’s performance on traffic speed datasets METR-LA, PEMS-BAY [31], and traffic flow dataset PEMS03 [32], each of which contains traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors. We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2. We predict the traffic speed/flow in the following 30 / 60 minutes (6 / 12 steps) from the previously observed 60 minutes (12 steps) for all datasets.
We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence. For undirected graphs ${\mathcal{G}}^{u}$, we connect each node with its $k$ nearest neighbors (NNs). We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a graph learning module before each ADMM block. We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism. We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths. See Appendix E for more details.
@@ 5.2 Experimental Results
Table 1: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03, METR-LA and PEMS-BAY datasets. We use boldface for the smallest error and underline the next 3 smallest errors.
Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges) PEMS-BAY (325 nodes, 2090 edges)
Horizon 30 minutes 60 minutes 30 minutes 60 minutes 30 minutes 60 minutes
VAR [33] 28.07 / 16.53 / 17.49 30.54 / 18.31 / 19.61 10.72 / 5.55 / 11.29 12.59 / 6.99 / 13.46 3.81 / 2.08 / 4.46 3.98 / 2.56 / 5.52
STGCN [34] 28.06 / 18.24 / 17.76 37.31 / 24.29 / 23.00 11.97 / 5.43 / 13.45 14.77 / 6.86 / 17.51 4.32 / 2.65 / 5.21 7.39 / 4.26 / 8.25
STSGCN [35] 26.41 / 16.75 / 16.86 30.62 / 19.35 / 19.15 10.25 / 4.05 / 9.18 12.65 / 5.18 / 11.48 3.43 / 1.54 / 3.28 4.50 / 1.93 / 4.42
ST-Wave [36] 25.57 / 15.11 / 15.04 28.65 / 16.81 / 19.24 10.81 / 4.11 / 9.16 13.24 / 5.33 / 11.32 3.53 / 1.49 / 3.18 4.30 / 1.82 / 4.33
ASTTN [9] 26.84 / 17.04 / 21.46 27.24 / 16.43 / 15.63 11.49 / 4.80 / 10.97 11.10 / 5.08 / 12.51 5.07 / 2.23 / 4.65 4.82 / 2.13 / 4.40
PDFormer [22] 23.71 / 15.05 / 18.16 27.16 / 17.26 / 21.21 10.21 / 3.89 / 8.50 12.17 / 4.81 / 10.92 3.18 / 1.41 / 2.99 4.55 / 2.00 / 4.88
STAEformer [37] 30.22 / 18.85 / 26.62 38.36 / 23.68 / 29.21 10.16 / 3.73 / 8.25 12.58 / 4.79 / 10.08 4.07 / 1.75 / 3.95 5.28 / 2.28 / 5.59
Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 10.35 / 4.22 / 9.70 12.37 / 5.27 / 11.91 3.78 / 1.70 / 4.08 4.60 / 2.03 / 4.79
@@
Table 2: Comparison of the number of parameters between our model and baseline deep learning models for 60-minute forecasting in PEMS03 dataset.
Model Ours STGCN STSGCN ST-Wave ASTTN PDFormer STAEformer
Params # 34,096 321,941 3,496,212 882,558 161,635 531,165 1,403,940
@@
We train each run of our model on NVIDIA GeForce RTX 3090. We evaluate our unrolling model against model-based method VAR [33], GNN-based models STGCN [^4] STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively. [34] and STSGCN [35], GAT-based models ST-Wave [36] and ASTTN [9], and transformer-based models PDFormer [22] and STAEformer [37] with 70 epochs of training on the reduced dataset with the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) in the predicted 6 / 12 steps.
@@
Table 1 shows the performance of our models and baselines in traffic forecasting, and Table 2 compares the number of parameters. We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer). Among all the models, ST-Wave generally achieves the best performance, which is attributed to an application-specific usage of wavelet transforms, but requires even more parameters than transformers. The largest model STSGCN fails to achieve satisfying performance for PEMS03, where the real connections are highly sparse. STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 and PEMS-BAY datasets. We also observe significant oscillations in validation loss during training in ASTTN due to the adaptive attention maps. In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the three datasets, while all require significantly more parameters than our models.
5.3 Ablation Study
Next, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations. First, we separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm. We denote these models as “w/o DGTV” and “w/o DGLR”. To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$. We denote the undirected-time unrolling model as “UT”. Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix F.
Dataset Mixed-Graph UT w/o DGTV w/o DGLR
METR-LA 12.37 / 5.27 / 11.91 12.51 / 5.42 /12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
For further ablation studies in the number of nearest neighbors $k$, size of time window $W$, number of ADMM blocks and layers, and the splitting-term strategy in (8), see Appendix G.

### v2

@@ 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in eq. 6 means
${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
We can compute eq. 15 element-wise as follows (see Appendix D for a derivation):
@@
Interpretation: Soft-thresholding in eq. 16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$. Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (Ortega et al., 2018), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli & Kovacevic, 2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$. Using eq. 16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved. For example, the constant (low-frequency) signal $\mathbf{1}$} passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$. Thus, attenuation of the high-pass channel means eq. 16 is a low-pass filter.
@@ 3.4.3 Updating Lagrange Multipliers
For objective eq. 6, we follow standard ADMM procedure to update Lagrange multiplier ${\bm{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
${\bm{\gamma}}^{\tau+1}$ $={\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\bm{\gamma}}_{u}^{\tau+1}$ and ${\bm{\gamma}}_{d}^{\tau+1}$ in objective eq. 9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~~~~~~~~~{\bm{\gamma}}_{d}^{\tau+1}={\bm{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
@@ 4 Unrolled Neural Network
We unroll our iterative ADMM algorithm to an interpretable neural net. The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (Bahdanau et al., 2014).
@@ 4.1 ADMM Algorithm Unrolling
We implement each iteration of our algorithm in Section 3.4 as a neural layer. Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems eq. 10 for ${\mathbf{x}}^{\tau+1}$, eq. 11 for ${\mathbf{z}}_{u}^{\tau+1}$, and eq. 12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via eq. 16. CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E for detailed implementation). Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer. Multipliers ${\bm{\gamma}}^{\tau+1},{\bm{\gamma}}_{u}^{\tau+1},{\bm{\gamma}}_{d}^{\tau+1}$ are then updated via eq. 17 and eq. 18 to complete one layer. Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\bm{\Theta}}_{b}$. See Fig. 3 for an illustration.
@@ 4.2 Self-Attention Operator in Transformer
We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (Bahdanau et al., 2014). Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$. The affinity between tokens $i$ and $j$, $e(i,j)$, is the dot product of linear-transformed ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{Q}},{\mathbf{K}}\in\mathbb{R}^{E\times E}$ are the query and key matrices, respectively. Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
@@ 4.3 Graph Learning Modules
Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$. For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng & Tassiulas (2022), and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details). As done in Thuc et al. (2024), the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
@@
where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$. Distance in eq. 21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$. Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
@@
Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in eq. 22 are essentially attention weights $a_{i,j}$’s in eq. 19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}. Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—eq. 10, eq. 11, eq. 12 and eq. 16—derived from our optimization of objective eq. 4.
Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$. For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$. Like eq. 21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}. Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization, as done in eq. 22 for an undirected edge.
Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\bm{\Phi}}_{b}$.
@@ 5.1 Experimental Setup
We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (Li et al., 2017) and traffic flow dataset PEMS03 (Guo et al., 2022), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors. We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2. We predict the traffic speed/flow in the following 30 / 60 /120 minutes (6 / 12 /24 steps) from the previously observed 60 minutes (12 steps) for all datasets. We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence. For undirected graphs ${\mathcal{G}}^{u}$, we connect each node to its $k$ nearest neighbors (NNs). We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a pair of graph learning modules before each ADMM block. We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism. We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths. See Appendix F for more details.
@@ 5.2 Experimental Results
Table 1: Comparison of categorized baselines and model parameters for 60-minute forecasting on the PEMS03 dataset against ours.
Category Selected Models Params #
Model-based VAR (Reinsel, 2003) -
[…]
GATs GMAN (Zheng et al., 2020), ST-Wave (Fang et al., 2023) 210K, 883K
Transformers PDFormer (Jiang et al., 2023), STAEformer (Liu et al., 2023) 531K, 1404K
Adaptive-graph models Graph WaveNet (GWN) (Wu et al., 2019), AGCRN (Bai et al., 2020) 277K, 749K
Simple linear models STID (Shao et al., 2022), SimpleTM (Chen et al., 2025) 123K, 540K
Mixed-graph unrolling Ours 34K
  • $\dagger$
    STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively.
Table 2: Comparison of RMSE / MAE / MAPE(%) metrics of our lightweight transformer to baseline models on 30 / 60-minute forecasting in PeMS03 and METR-LA datasets. We use boldface for the smallest error, color the 2nd and 3nd small error in blue, and underline the next 2 smallest errors.
Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges)
Horizon 30 minutes 60 minutes 120 minutes 30 minutes 60 minutes 120 minutes
VAR 28.07 / 16.53 / 17.49 30.54 / 18.31 / 19.61 36.65 / 22.40 / 24.64 10.72 / 5.55 / 11.29 12.59 / 6.99 / 13.46 14.83 / 8.90 / 16.54
[…]
STSGCN 26.41 / 16.75 / 16.86 30.62 / 19.35 / 19.15 38.04 / 23.86 / 23.62 10.25 / 4.05 / 9.18 12.65 / 5.18 / 11.48 15.74 / 6.84 / 15.33
[…]
ST-Wave 25.57 / 15.11 / 15.04 28.65 / 16.81 / 19.24 29.88 / 17.11 / 17.71 10.81 / 4.11 / 9.16 13.24 / 5.33 / 11.32 23.18 / 11.22 / 12.71
[…]
STAEformer 30.22 / 18.85 / 26.62 38.36 / 23.68 / 29.21 48.72 / 31.96 / 44.71 10.16 / 3.73 / 8.25 12.58 / 4.79 / 10.08 14.63 / 5.82 / 11.87
[…]
AGCRN 27.40 / 15.19 / 14.35 29.90 / 16.76 / 15.32 32.79 / 18.72 / 17.03 10.11 / 3.82 / 8.61 12.56 / 5.00 / 11.25 14.77 / 6.33 / 14.05
[…]
SimpleTM 23.75 / 15.13 / 15.59 25.56 / 15.97 / 15.49 30.58 / 18.73 / 18.18 8.76 / 4.12 / 7.63 11.96 / 5.49 / 9.65 15.44 / 7.64 / 12.27
Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 34.40 / 21.16 / 23.47 10.35 / 4.22 / 9.70 12.34 / 5.18 / 11.80 15.19 / 7.56 / 16.12
@@
We train each run of our model on an entire NVIDIA GeForce RTX 3090. We evaluate our unrolling model against the baselines shown in Table 1. Each model is trained for 70 epochs on the reduced dataset, and the metrics are the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) in the predicted 6 / 12 / 24 steps.
@@
Table 2 shows the performance of our models and baselines in traffic forecasting, and the last column of Table 1 compares the number of parameters. We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer). Among all the models, ST-Wave and SimpleTM generally achieve the best performance. However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (Battaglia et al., 2018); both models require even more parameters than PDFormer. Our unrolled model also generally outperforms the multi-attention model GMAN. Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance. The largest GCN model, STSGCN, fails to achieve a satisfying performance for PEMS03, where the real connections are highly sparse. STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 datasets. In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the two datasets, while all require significantly more parameters than our models. In contrast, our model captures the complex and evolving spatial-temporal relationship accurately using a mixed graph, demonstrating consistent performance for all predicting horizons; notably, the performance gap to the best model decreases as the horizon grows.
Further experiments on data efficiency are presented in Appendix G, and a comprehensive ablation study is provided in Appendix H.